can a 34 inch baseball bat fit into a 24in by 20in by 12in box?

Answer by user_dude2008(1862)

sqrt(24^2+20^2+12^2)=sqrt(576+400+144)=sqrt(1120)=33.4664011

Answer: no it cannot.